OpenStudy (anonymous):

Find x if y = –2. 4x-6/y=15

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yup thanks

OpenStudy (anonymous):

Wouldn't x= 6 because you divided the 4x-6 by 2 to get -2x+3=15. then subtract -3 and divide by -2. Check it

OpenStudy (anonymous):

No your doing distribution. If you go step by step youll under stand how my answers right. Heres the work. Look and see. Find x if y = –2. 4x-6/-2=15 4x+3=15 4x=12 x=3

OpenStudy (anonymous):

You want me to explain it?

OpenStudy (anonymous):

no thanks